In the month of March, Bramble Salon services 550 clients at an average price of $130. During the month, fixed costs were $23Using the contribution margin technique, compute the break-even point in dollars and in units. Break-even sales Break-even saj
Answer #1

1. Contribution Margin = Sales - Variable cost

Sales = 550 clients * $130 per client

= $ 71,500

Variable cost = 70% of sales i.e. $ 50,050

Thus, Contribution margin = 71500 - 50050 = $ 21,450

2. Contribution margin per unit = Contribution margin/ no. of units

= $ 21,450/ 550 clients

= $ 39 per unit

3. Contribution margin ratio = 30%

Variable costs are 70% of sales, Hence, contribution ratio = (100 - 70) i.e. 30%

4. Break even sales (in units) = Total fixed cost/ Contribution per unit

= $ 23,400/ 39

= 600 clients

5. Break even sales (in $) = 600 clients * $ 130 per client

= $ 78,000

In the given scenario, as the company is unable to achieve it's break even i.e. 600 clients, it has to increase it's production level, as currently the company is incurring losses.